Write the first five terms of the sequence.
a n = n^2 + 2

Answer by greenestamps(13200)



The definition means that the value of the n-th term is n squared plus 2.

The 1st term is then 1^2+2 = 1+2 = 3.
The 2nd term is 2^2+2 = 4+2 = 6.
I think you can find the 3rd, 4th, and 5th terms.